# Path integral and conformal instability in nonlocal quantum gravity

### v1

@@ Abstract
We introduce the Lorentzian path integral of nonlocal quantum gravity. After introducing the functional measure, the Faddeev–Popov sector, and the field correlators, we move to perturbation theory and describe Efimov analytic continuation of scattering amplitudes to Euclidean momenta and back to Lorentzian. We show that the conformal instability problem in the Euclidean path integral is solved by suitable gauge choices at the perturbative level. The four examples of Einstein gravity, Stelle gravity, minimally coupled nonlocal quantum gravity, and nonminimally coupled nonlocal quantum gravity are given.
@@
Models of Quantum Gravity, Gauge Symmetry
@@ 1 Introduction
In the last five decades, the quest for a theory where the gravitational interaction and quantum mechanics can be combined consistently has seen a resurgence of candidates BMS , to the point where it is no longer tenable to claim that we do not know how to quantize gravity. If anything, the problem is that we know far too many ways to do it and too few ways to test their predictions with observations. Supergravity WeB ; FvP , the low-energy limit of string theory GSWi ; Pol98 ; BBSb ; Zwi09 , group field theory Ori09 ; BaO11 ; Fousp , loop quantum gravity and spin foams Rov07 ; Per13 ; Ashtekar:2021kfp ; Engle:2023qsu , asymptotically safe quantum gravity Niedermaier:2006wt ; Eichhorn:2018yfc ; Bonanno:2020bil ; Saueressig:2023irs , causal dynamical triangulations Ambjorn:2012jv ; Loll:2019rdj ; Ambjorn:2024pyv and others combine different notions of spacetime, quantization methods and dynamics. The degree of success in quantizing gravity and the level of completion of our understanding of the properties and phenomenology of the theory greatly vary among these scenarios. While some of them are nonperturbative, others are based on perturbative quantum field theory (QFT), let it be defined on a group manifold as in group field theory, on a higher-dimensional smooth spacetime where all fields enjoy supersymmetry as in supergravity, or on more minimalistic scenarios with a four-dimensional spacetime and without supersymmetry. In the latter case, of course, the price to pay to stay in a conservative perturbative QFT setting is to add or modify some of the traditional ingredients, such as making the dynamics fundamentally nonlocal (nonlocal quantum gravity, minimally Kra87 ; Kuzmin:1989sp ; Tomboulis:1997gg ; Modesto:2011kw ; Biswas:2011ar ; BCKM ; Modesto:2014lga ; Dona:2015tra ; Koshelev:2016xqb ; Modesto:2016max ; CMN3 ; Modesto:2017sdr ; Briscese:2018oyx ; Buoninfante:2018mre ; BasiBeneito:2022wux ; Buoninfante:2022ild or nonminimally Modesto:2021ief ; Modesto:2021okr ; Modesto:2021soh ; Modesto:2022asj ; Calcagni:2023goc coupled to matter), imposing a prescription on the propagators and a projection on the spectrum to get rid of physical ghost modes (fakeon gravity Anselmi:2017yux ; Anselmi:2017lia ; Anselmi:2018kgz ; Anselmi:2018ibi ; Anselmi:2018tmf ; Anselmi:2018bra ; Anselmi:2019rxg ; Anselmi:2019nie ; Anselmi:2021hab ), or giving up the notion of point particles in favour of a gas of quasi-particles living on a fractal spacetime (fractional gravity Calcagni:2016azd ; Calcagni:2021ipd ; Calcagni:2021aap ; Calcagni:2022shb ).
@@
Among the most recent perturbative approaches, the one has perhaps received more attention is nonlocal quantum gravity. The action of the theory is characterized by asymptotically polynomial nonlocal operators, entire functions of the Laplace–Beltrami operator $\Box$ which do not add extra poles in the propagator while, at the same time, taking the form of finite-order polynomials in the ultraviolet (UV). Thanks to these properties, perturbative unitarity is preserved, a power-counting analysis of divergences is possible and there are robust indications that the theory is super-renormalizable or, in some of its versions, finite. Despite these advances, however, certain basic aspects of the theory have not been discussed with the due emphasis. For example, many works have been devoted to Feynman diagrams and scattering amplitudes but little has been said about their origin from a path integral. Moreover, these diagrams and amplitudes are invariably calculated in Euclidean momentum space due to the difficulty, or even impossibility, to handle a well-defined nonlocal QFT exclusively in Lorentzian signature. Therefore, questions may arise on whether the fundamental formulation of nonlocal quantum gravity is based on a Lorentzian or a Euclidean path integral, whether such path integral is convergent, and so on.
@@
where $\mathcal{L}$ is the Lagrangian, we work in mostly plus signature (,+,.+)formulae-sequence(-,+$\cdots$.+)( - , + , ⋯ . + ) and $M_{\text{\tiny Pl}}\coloneqq(8\pi G)^{-1/2}$ is the reduced Planck mass with energy dimensionality $[M_{\text{\tiny Pl}}]=(D-2)/2$. When giving a concrete name to the Lagrangian, we will consider four theories:
@@
    where ${\rm H}_{0,2}(\Box)$ are two entire functions which, in this class of theories, are asymptotically polynomial in the ultraviolet (UV) and such that ${\rm H}_{0,2}(0)=0$ in the infrared (IR). The reason for this specific form of nonlocal operators will be discussed in section 3. For simplicity, one can set ${\rm H}_{0}(\Box)={\rm H}_{2}(\Box)={\rm H}(\Box)$ but we will not do that until later. Due to the Gauss–Bonnet theorem, the Riemann–Riemann term and its associated form factor $\gamma_{4}$ play no role in the renormalizability of the theory, since its presence amounts to a redefinition of the form factors appearing in the graviton propagator and scattering amplitudes:
@@
    Then, the action reduces to the same as in the minimally coupled case, with $\gamma_{4}=0$. The form factors are of asymptotically polynomial form,
@@
In section 2.5, we recall the analytic continuation of scattering amplitudes to Euclidean signature. In section 3, we calculate the free-level (e.g., tree-level) graviton propagator for the generic theory (4) in a general gauge, without specifying the form factors; the final result applies to all the above cases upon the choice of $\gamma_{0,2,4}$. In section 4, we discuss the conformal instability problem in the Euclidean path integral and show that it can be solved by gauge invariance at the perturbative level for any of the above theories. In general, the conformal instability arises when one ignores gauge-fixing terms in the action, which are necessary to invert the graviton kinetic term. Section 5 contains our conclusions and future extensions of these results, in particular, at the nonperturbative level.
@@ 2 Lorentzian path integral
The path-integral quantization of nonlocal quantum gravity in the background-field gauge follows the same line as that of higher-derivative local theories BOS . In this section, we consider only the gravitational sector of nonlocal quantum gravity, i.e., the action (4). Introduce the background metric $\tilde{g}_{\mu\nu}$ and the Green’s functions generating functional
@@
where $\chi_{\alpha}$ and the matrix $M^{\alpha}_{\beta}$ depend both on the total metric $g_{\mu\nu}$ and on the background metric $\tilde{g}_{\mu\nu}$ separately. The gauge condition is expressed in terms of a functional $t^{\alpha}_{\mu\nu}[\tilde{g}]$ depending only on the background metric $\tilde{g}_{\mu\nu}$,
@@ 2.1 Quantum effective action
$\Gamma[\tilde{g},\bar{g}]=W[\tilde{g},J]-\int_{\tilde{g}}\bar{g}_{\mu\nu}J^{\mu\nu}.$ (23)
@@
where the last integral is $\int_{\eta}\coloneqq\int d^{D}x$ because the $1/\sqrt{|\tilde{g}|}$ prefactor in eq. (24) cancels the one in the integral measure.
@@
$t_{\alpha}^{\mu\nu}[\tilde{g}]=\tilde{g}^{\mu\sigma}\delta^{\nu}_{\alpha}\,\tilde{\nabla}_{\sigma}-\beta_{g}\,\tilde{g}^{\mu\nu}\tilde{\nabla}_{\alpha}\,,$ (33)
@@
$\chi_{\alpha}[\tilde{g},h]=t_{\alpha}^{\mu\nu}[\tilde{g}]\,h_{\mu\nu}=\left(\tilde{g}^{\mu\sigma}\delta^{\nu}_{\alpha}\,\tilde{\nabla}_{\sigma}-\beta_{g}\,\tilde{g}^{\mu\nu}\tilde{\nabla}_{\alpha}\right)h_{\mu\nu}={\tilde{\nabla}}_{\sigma}h^{\sigma}_{\alpha}-\beta_{g}{\tilde{\nabla}}_{\alpha}h\,.$ (34)
@@
$=$ ~${\tilde{\Box}\delta^{\alpha}_{\beta}+\tilde{\nabla}_{\beta}\tilde{\nabla}^{\alpha}-2\beta_{g}\tilde{\nabla}^{\alpha}\tilde{\nabla}_{\beta}}+\tilde{\nabla}^{\mu}\left(h_{\mu\beta}\tilde{\nabla}^{\alpha}\right)$
+~$+\tilde{\nabla}^{\mu}\left(h^{\alpha}_{\beta}\tilde{\nabla}_{\mu}\right)-2\beta_{g}\tilde{\nabla}^{\alpha}\left(h_{\mu\beta}\tilde{\nabla}^{\mu}\right),$
@@
where $\gamma_{2}^{\prime}=\gamma_{2}^{\prime}(\tilde{\Box})$ and the gauge-fixing parameters $\lambda$, $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are dimensionless constants, $[\lambda]=[\lambda_{i}]=0$, just like for a local theory. [^1] As we will see in section 4, there is no need to generalize the $\lambda$ and $\lambda_{i}$ to nontrivial operators. In eq. (38), we used the compatibility condition ~$\tilde{\nabla}_{\alpha}\,\tilde{g}_{\beta\gamma}=0$ for the background metric. Note that, on a curved background, covariant derivatives do not commute, hence the separation of the last two factors in (38). The expression for $\mathcal{G}^{\mu\nu}$ generalizes the one of Asorey:1996hz to the case of nonlocal theories.
@@ 2.3 One-loop effective action
In order to derive the one-loop quantum effective action we have to expand the action at the second order in the fields. In particular, we expand the gravitational action at the second order in $h_{\mu\nu}$, so that no gravitational coupling appears between the ghosts and the graviton. Therefore, (29) simplifies to
@@ 2.4 Green’s functions
This subsection is devoted to the path-integral formulation of any gravitational theory in Minkowski spacetime. Namely, the background spacetime is globally Minkowski. Here we treat gravity on exactly the same status as the other fundamental interactions in Nature. Indeed, the choice of the Minkowski background in gravity is analogue to the trivial zero background for gauge bosons, scalars, or fermions and it is dictated by the equivalence principle.
@@ 2.5 Efimov analytic continuation
It has become progressively clear that even the best behaved nonlocal quantum field theories have problems if scattering amplitudes are calculated directly in Lorentzian signature Briscese:2021mob . If, instead, one defines the Feynman diagrams with imaginary external and internal energies (Euclidean signature) and then analytically continues external momenta to real energies after integrating, then loop integrals can be performed consistently and the theory admits a unique Lorentzian limit. This is the so-called Efimov analytic continuation Efimov:1967dpd ; Pius:2016jsl ; Briscese:2018oyx , which consists of three steps. For any given Feynman diagram in Lorentzian quantum gravity:
@@
    Define a Euclidean version $k_{4}=ik^{0}$ and $p_{4}=ip^{0}$ of both internal ($k$) and external ($p$) energies, where $k^{0}$ and $p^{0}$ are the real energies of the original theory.
@@
    Analytically continue both internal and external energies to the complex plane. In this way, integration over $k_{4}$ is performed on an open path on the imaginary axis. (Note that, in general, there will be poles $\bar{k}_{4}(p)$ of the integrand to the left and to the right of the path, but none on the path itself Briscese:2018oyx .)
@@
    Send the external energies $p_{4}$ back to purely imaginary (real $p^{0}$) and deform the path in such a way as to accommodate the poles that migrate across the imaginary axis. This guarantees safe integration along the path.
At the perturbative level, there is no conceptual difference between gravity and any other interactive gauge theory. Perturbative quantum gravity is the theory of interacting gravitons and loop amplitudes can be computed exactly as in those cases. All loop amplitudes are computed with purely imaginary internal and external energies and, in the end, one makes the above analytic continuation and the final result is a Lorentzian quantum effective action. The only difference with respect to two-derivative theories is that the analytically continued theory is not equivalent to the theory defined with internal real energies because of the contribution of the pole at infinity (essential singularity) due to the nonlocal form factor.

### v2

@@ Abstract
We introduce the Lorentzian path integral of nonlocal quantum gravity. After introducing the functional measure, the Faddeev–Popov sector and the field correlators, we move to perturbation theory and describe Efimov analytic continuation of scattering amplitudes to Euclidean momenta and back to Lorentzian. We show that the conformal instability problem in the Euclidean path integral is solved by suitable gauge choices at the perturbative level. The three examples of Einstein gravity, Stelle gravity and nonlocal quantum gravity are given.
@@
Gauge Symmetry, Models of Quantum Gravity
@@ 1 Introduction
In the last five decades, the quest for a theory where the gravitational interaction and quantum mechanics can be combined consistently has seen a resurgence of candidates BMS , to the point where, although we do not yet have a single fully consistent and controllable theory dominating the others, it is no longer tenable to claim that we do not know at all how to quantize gravity. If anything, the problem is that we are aware of far too many ways in which we could do it and too few ways to test their predictions with observations. Supergravity WeB ; FvP , the low-energy limit of string theory GSWi ; Pol98 ; BBSb ; Zwi09 , group field theory Ori09 ; BaO11 ; Fousp , loop quantum gravity and spin foams Rov07 ; Per13 ; Ashtekar:2021kfp ; Engle:2023qsu , asymptotically safe quantum gravity Niedermaier:2006wt ; Eichhorn:2018yfc ; Bonanno:2020bil ; Saueressig:2023irs , causal dynamical triangulations Ambjorn:2012jv ; Loll:2019rdj ; Ambjorn:2024pyv and others combine different notions of spacetime, quantization methods and dynamics. The degree of success in quantizing gravity and the level of completion of our understanding of the properties and phenomenology of the theory greatly vary among these scenarios. While some of them are nonperturbative, others are based on perturbative quantum field theory (QFT), let it be defined on a group manifold as in group field theory, on a higher-dimensional smooth spacetime where all fields enjoy supersymmetry as in supergravity, or on more minimalistic scenarios with a four-dimensional spacetime and without supersymmetry. In the latter case, of course, the price to pay to stay in a conservative perturbative QFT setting is to add new ingredients or to modify some of the traditional ones, such as making the dynamics fundamentally nonlocal (nonlocal quantum gravity, minimally Kra87 ; Kuzmin:1989sp ; Tomboulis:1997gg ; Modesto:2011kw ; Biswas:2011ar ; BCKM ; Modesto:2014lga ; Dona:2015tra ; Koshelev:2016xqb ; Modesto:2016max ; CMN3 ; Modesto:2017sdr ; Briscese:2018oyx ; Buoninfante:2018mre ; BasiBeneito:2022wux ; Buoninfante:2022ild or nonminimally Modesto:2021ief ; Modesto:2021okr ; Modesto:2021soh ; Modesto:2022asj ; Calcagni:2023goc coupled to matter), imposing a prescription on the propagators and a projection on the spectrum to get rid of physical ghost modes (fakeon gravity Anselmi:2017yux ; Anselmi:2017lia ; Anselmi:2018kgz ; Anselmi:2018ibi ; Anselmi:2018tmf ; Anselmi:2018bra ; Anselmi:2019rxg ; Anselmi:2019nie ; Anselmi:2021hab ), or giving up the notion of point particles in favour of a gas of quasi-particles living on a fractal spacetime (fractional gravity Calcagni:2016azd ; Calcagni:2021ipd ; Calcagni:2021aap ; Calcagni:2022shb ).
@@
Among the most recent perturbative approaches, the one that has perhaps received more attention is nonlocal quantum gravity or, more precisely, asymptotically local quantum gravity. The action of the theory is characterized by asymptotically polynomial nonlocal operators, entire functions of the Laplace–Beltrami operator $\Box$ which do not add extra poles in the propagator while, at the same time, taking the form of finite-order polynomials in the ultraviolet (UV). Thanks to these properties, perturbative unitarity is preserved, a power-counting analysis of divergences is possible and there are robust indications that the theory is super-renormalizable or, in some of its versions, finite. Despite these advances, however, certain basic aspects of the theory have not been discussed with the due emphasis. For example, many works have been devoted to Feynman diagrams and scattering amplitudes but little has been said about their origin from a path integral. Moreover, these diagrams and amplitudes are invariably calculated in Euclidean momentum space due to the difficulty, or even impossibility, to handle a well-defined nonlocal QFT exclusively in Lorentzian signature. Therefore, questions may arise on whether the fundamental formulation of nonlocal quantum gravity is based on a Lorentzian or a Euclidean path integral, whether such path integral is convergent, and so on.
@@
where $\mathcal{L}$ is the Lagrangian, we work in mostly plus signature (,+,,+)(-,+$\cdots$,+)( - , + , ⋯ , + ) and $M_{\text{\tiny Pl}}\coloneqq(8\pi G)^{-1/2}$ is the reduced Planck mass with energy dimensionality $[M_{\text{\tiny Pl}}]=(D-2)/2$. When giving a concrete name to the Lagrangian, we will consider four theories:
@@
    where ${\rm H}_{0,2}(\Box)$ are two entire functions which, in the case of asymptotically local quantum gravity, are asymptotically polynomial in the ultraviolet (UV) and such that ${\rm H}_{0,2}(0)=0$ in the infrared (IR). The reason for this specific form of nonlocal operators will be discussed in section 3. For simplicity, one can set ${\rm H}_{0}(\Box)={\rm H}_{2}(\Box)={\rm H}(\Box)$ but we will not do that until later. Due to the Gauss–Bonnet theorem, the Riemann–Riemann term and its associated form factor $\gamma_{4}$ play no role in the renormalizability of the theory, since its presence amounts to a redefinition of the form factors appearing in the graviton propagator and scattering amplitudes:
@@
    Then, the action reduces to the same as in the minimally coupled case, with $\gamma_{4}=0$. The form factors in the class of nonlocal theories considered here are of asymptotically polynomial type,
@@
In section 2.5, we recall the analytic continuation of scattering amplitudes to Euclidean signature. In section 3, we calculate the free-level (i.e., tree-level) graviton propagator for the generic theory (4) in a general gauge, without specifying the form factors; the final result applies to all the above cases upon the choice of $\gamma_{0,2,4}$. In section 4, we discuss the conformal instability problem in the Euclidean path integral and show that it can be solved by gauge invariance at the perturbative level for any of the above theories. In general, the conformal instability arises when one ignores gauge-fixing terms in the action, which are necessary to invert the graviton kinetic term. Section 5 contains our conclusions and future extensions of these results, in particular, at the nonperturbative level.
@@ 2 Lorentzian path integral
The path-integral quantization of nonlocal quantum gravity in the background-field gauge follows the same line as that of higher-derivative local theories BOS . In this section, we consider only the gravitational sector of nonlocal quantum gravity, i.e., the action (4), without specifying the type of form factors (hence, what follows applies also to fractional gravity Calcagni:2016azd ; Calcagni:2021ipd ; Calcagni:2021aap ; Calcagni:2022shb ). Introduce the background metric $\tilde{g}_{\mu\nu}$ and the Green’s functions generating functional
@@
where $\chi_{\alpha}$ and the matrix $M^{\alpha}_{\beta}$ depend both on the total metric $g_{\mu\nu}$ and on the background metric $\tilde{g}_{\mu\nu}$ separately. The gauge condition is expressed in terms of a functional $t_{\alpha}^{\mu\nu}[\tilde{g}]$ depending only on the background metric $\tilde{g}_{\mu\nu}$,
@@ 2.1 Quantum effective action
$\Gamma[\tilde{g},\bar{g}]\coloneqq W[\tilde{g},J]-\int_{\tilde{g}}\bar{g}_{\mu\nu}J^{\mu\nu}.$ (23)
@@
where the first spacetime integral is $\int_{\eta}\coloneqq\int d^{D}x$ because the $1/\sqrt{|\tilde{g}|}$ prefactor in eq. (24) cancels the one in the integral measure.
@@
$t_{\alpha}^{\mu\nu}[\tilde{g}]=\tilde{g}^{\mu\sigma}\delta^{\nu}_{\alpha}\,\tilde{\nabla}_{\sigma}-\lambda\,\tilde{g}^{\mu\nu}\tilde{\nabla}_{\alpha}\,,$ (33)
@@
$\chi_{\alpha}[\tilde{g},h]=t_{\alpha}^{\mu\nu}[\tilde{g}]\,h_{\mu\nu}=\left(\tilde{g}^{\mu\sigma}\delta^{\nu}_{\alpha}\,\tilde{\nabla}_{\sigma}-\lambda\,\tilde{g}^{\mu\nu}\tilde{\nabla}_{\alpha}\right)h_{\mu\nu}={\tilde{\nabla}}_{\sigma}h^{\sigma}_{\alpha}-\lambda{\tilde{\nabla}}_{\alpha}h\,.$ (34)
@@
$=$ ~${\tilde{\Box}\delta^{\alpha}_{\beta}+\tilde{\nabla}_{\beta}\tilde{\nabla}^{\alpha}-2\lambda\tilde{\nabla}^{\alpha}\tilde{\nabla}_{\beta}}+\tilde{\nabla}^{\mu}\left(h_{\mu\beta}\tilde{\nabla}^{\alpha}\right)$
+~$+\tilde{\nabla}^{\mu}\left(h^{\alpha}_{\beta}\tilde{\nabla}_{\mu}\right)-2\lambda\tilde{\nabla}^{\alpha}\left(h_{\mu\beta}\tilde{\nabla}^{\mu}\right),$
@@
where $\gamma_{2}^{\prime}=\gamma_{2}^{\prime}(\tilde{\Box})$ and the gauge-fixing parameters $\lambda$, $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are dimensionless constants, $[\lambda]=[\lambda_{i}]=0$, just like for a local theory. [^1] As we will see in section 4, there is no need to generalize $\lambda$ and $\lambda_{i}$ to non-trivial operators. In eq. (38), we used the compatibility condition ~$\tilde{\nabla}_{\alpha}\,\tilde{g}_{\beta\gamma}=0$ for the background metric. Note that, on a curved background, covariant derivatives do not commute, hence the separation of the last two factors in (38). The expression for $\mathcal{G}^{\mu\nu}$ generalizes the one of Asorey:1996hz to the case of nonlocal theories.
@@ 2.3 One-loop effective action
In order to derive the one-loop quantum effective action, we have to expand the action at the second order in the fields. In particular, we expand the gravitational action at the second order in $h_{\mu\nu}$, so that no gravitational coupling appears between the ghosts and the graviton. Therefore, (29) simplifies to
@@
The effective action in nonlocal quantum gravity has not been calculated in full yet and super-renormalizability has been checked so far with the power-counting argument and via the calculation of the one-loop effective action in scalar toy models and in gravitational models with fewer derivatives than those appearing in the UV limit of asymptotically polynomial form factors, in particular, six Rachwal:2021bgb . However, increasing the number of derivatives makes the power counting even more powerful because the beta functions do not depend on the running couplings. In general, divergences come only from the polynomial (local) parts of the theory, while diagrams with nonlocal form factors within are, by definition, all convergent on the domain of such operators. It is important to stress that power counting is valid, as a general argument, to conclude that the theory is super-renormalizable. The reason is that the counter-terms to be added to the bare Lagrangian are local operators Calcagni:2023goc , which implies that the Bogoliubov–Parasiuk–Hepp–Zimmermann (BPHZ) subtraction scheme holds and that the theory is also BPHZ renormalizable Calcagni:2022shb . This conclusion is somewhat obvious from the fact that the UV limit of nonlocal theories with asymptotically polynomial form factors is local but, surprisingly, it also holds for some nonlocal theories with a nonlocal UV limit such as fractional gravity Calcagni:2022shb . An immediate consequence is that BPHZ renormalization guarantees that sub-divergences can always be absorbed by a standard one-loop counter-term in the action. Hence, the power-counting argument is sufficient to account for all divergences at any loop order. Note also that these complications do not appear in the finite version of the theory. In this case, there are no sub-divergences at higher loop orders because there are no divergences at all at lower order, since all beta functions vanish Modesto:2014lga ; Calcagni:2023goc . Thanks to these general results, the explicit calculation of the one-loop quantum effective action (48) becomes secondary in the discussion of the renormalizability of nonlocal quantum gravity, although, of course, it should be on top of the future agenda.
@@ 2.4 Green’s functions
This subsection is devoted to the path-integral formulation of any gravitational theory in Minkowski spacetime. Namely, the background spacetime is globally Minkowski. Here we treat gravity in exactly the same way as the other fundamental interactions in Nature. Indeed, the choice of the Minkowski background in gravity is analogue to the trivial zero background for gauge bosons, scalars, or fermions and it is dictated by the equivalence principle.
@@ 2.5 Efimov analytic continuation
It has become progressively clear that even the best behaved nonlocal quantum field theories have problems if scattering amplitudes are calculated directly in Lorentzian signature Briscese:2021mob . In other words, usually, naive Lorentzian nonlocal theories where the internal energies $k^{0}$ are integrated on the real line do not exist. If, instead, one defines the Feynman diagrams with imaginary external and internal energies (Euclidean signature in momentum space) and afterwards analytically continues external energies to real ones after integrating, then loop integrals can be performed consistently and the theory admits a unique Lorentzian limit. This is the so-called Efimov analytic continuation Efimov:1967dpd ; Pius:2016jsl ; Briscese:2018oyx ; Chin:2018puw ; Koshelev:2021orf ; Buoninfante:2022krn , which consists of three steps.
For any given loop Feynman diagram in Lorentzian quantum gravity, namely, where the path integral is performed on Lorentzian metrics (perturbations on Minkowski spacetime) with the $i\epsilon$ Feynman prescription:
@@
    Assume that both internal and external energies (respectively, $k^{0}$ and $p^{0}$) take complex values in the loop amplitudes.
@@
    Integrate on the imaginary axis in the $(\text{Re}\,k^{0},\,\text{Im}\,k^{0})$ complex plane. (Note that, in general, there will be poles $\bar{k}^{0}(p)$ of the integrand to the left and to the right of the path, but none on the path itself Briscese:2018oyx .)
@@
    After the loop integrations, analytically continue the external energies $p^{0}$ back to real values.
We can express the above three steps into other, equivalent ways.
One, which we might call deformed-paths view, is to assume that only the internal energies $k^{0}$ are complex valued, while the external energies $p^{0}$ stay real. Then, internal energies are integrated along special paths in the complex plane. According to the previous description, these paths are obtained moving analytically the external energies from complex to real values. Each path is deformed in such a way as to accommodate analytically the poles that migrate across the imaginary $k^{0}$ axis, thus guaranteeing safe integration Briscese:2018oyx . The advantage of understanding Efimov analytic continuation according to steps 1-3 instead of in the deformed-paths view is that the correct deformed path and the homotopic class it represents can hardly be guessed if external energies stay real and internal ones are imaginary.
Another equivalent form (Euclidean theory view) is to invoke the Euclidean version of the theory, which corresponds to the third line of (LABEL:scheme) below. In step 2, define first a Euclidean version $k_{D}=-ik^{0}$ and $p_{D}=-ip^{0}$ of both internal and external energies and integrate on real values of $k_{D}$. Obviously, in Euclidean quantum gravity this step would be automatically implemented. Then, it is clear that the Lorentzian amplitudes of the nonlocal theory are nothing but Euclidean amplitudes analytically continued to imaginary external energies $p_{D}$. Therefore, at the perturbative level, the Lorentzian theory as defined in sections 2.12.4 is indistinguishable from an analytically continued Euclidean version of the theory.
Of course, this means that the Lorentzian and the pure Euclidean theory (i.e., without analytic continuation) are physically inequivalent, as also noted in Buoninfante:2022krn , since there is a highly non-trivial analytic continuation differentiating the two. In the pure Euclidean case, one integrates straight on top of the real axis in the $(\text{Re}\,k_{D},\text{Im}\,k_{D})$ plane, and this is it. In contrast, in the Lorentzian case one deforms this path when sending the external energies to real values and some poles get across the imaginary axis Briscese:2018oyx . This is also the reason why we started with a Lorentzian path-integral formulation instead of a Euclidean one: the final result is unique because the power counting and all the integrals are defined in Euclidean momentum space.
Summarizing, at the perturbative level there is no conceptual difference between gravity and any other interactive gauge theory. Perturbative quantum gravity is the theory of interacting gravitons and loop amplitudes can be computed exactly as in those cases. All loop amplitudes are computed with purely imaginary internal and external energies and, in the end, one makes the above analytic continuation and the final result is a Lorentzian quantum effective action. The only difference with respect to two-derivative theories is that the analytically continued theory is not equivalent to the theory defined with internal real energies because of the contribution of the pole at infinity (essential singularity) due to the nonlocal form factor.